A conducting bar moves along frictionless conducting rails connected to a 4.00 omega resistor. The length of the bar is 1.60m and a uniform magnetic field of 2.20T is applied perpendicular to the paper pointing outward as shown

a) What is the applied force required to move the bar to the right with a constant speed of 6.00 m/s?

b) At what rate is energy dissipated in the 4.00 ohm resistor?
Answer #1

E= Blv

=2.20 x 1.6 x 6

=21.12V

now I = E/R = 21.12/4 = 5.28A

a. force on the rod = IlB = 5.28 x 1.6 x 2.2 = 18.58 N

b. rate of dissipation of energy = I^2R =5.28 x 5.28 x 4 =111.5 W